OpenStudy (anonymous):

Solve 3(4x - 3) = 51

OpenStudy (lilg132):

12x - 9 = 51

OpenStudy (lilg132):

12x = 51 + 9 12x = 60 12x / 12 = 60 / 12 x = 5

OpenStudy (anonymous):

thankyou! :)can you help me with this too? Evaluate the following expression using the values given. Find 4x - y - 2z if x = -2, y = 3, and z = -2

OpenStudy (lilg132):

click good answer?

OpenStudy (lilg132):

4(-2) -(3) -2(-2)